OpenStudy (anonymous):

please help. 9x^2-42xy+49y^2 will medal and fan if helped

OpenStudy (anonymous):

factor completely

OpenStudy (fibonaccichick666):

so what do you know about the numbers 9 and 49?

DivineSolar (divinesolar):

Does prime happen to be one of your answers! :P

OpenStudy (anonymous):

they have a nine in them jk umm... idk

OpenStudy (anonymous):

nope @DivineSolar

DivineSolar (divinesolar):

Well here. 9x^2 - 42xy + 49y^2 = (3x)^2 - 42xy + (7y)^2 = (3x)^2 - 2 * (3x) * (7y) + (7y)^2 It is in the form of a^2 - 2ab +b^2 i.e, (a-b)^2 = (3x - 7y)^2

OpenStudy (anonymous):

it is similar to the other trinomial posted earlier

OpenStudy (anonymous):

yeah true @clickspiker23

DivineSolar (divinesolar):

Is my final answer an answer for your question conner?

OpenStudy (fibonaccichick666):

@DivineSolar please remember that we are here to guide askers to solutions-not just answer it for them. http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

looks like @DivineSolar is ahead of me

DivineSolar (divinesolar):

Well fibonacci i explained it by working it out. And you also are not a moderator ^-^

OpenStudy (fibonaccichick666):

So anyways, have you ever heard of a perfect square connor?

DivineSolar (divinesolar):

And yes i understand you are just giving out a fair warning ^-^

OpenStudy (anonymous):

is that the yes it is thank you

OpenStudy (fibonaccichick666):

@DivineSolar that is not following the code of conduct as it is not guiding out to a solution. 2. in order to prevent such things occurring it is every member's duty to uphold the code and inform those infringing upon it.

DivineSolar (divinesolar):

Yes but you are breaking the COC by spamming a topic unrelated to the question

OpenStudy (fibonaccichick666):

I am sorry. @conner143 I do not understand this statement "is that the yes it is thank you " can you please elaborate?

OpenStudy (anonymous):

that was the answer to the question

DivineSolar (divinesolar):

More than once counts as spam. Once its said its final.

OpenStudy (fibonaccichick666):

Well, you should be able to figure out if it is correct by Foiling.

OpenStudy (fibonaccichick666):

@DivineSolar if you feel you can report that and it be effective, please do. However, you still fail to understand what I was saying.

DivineSolar (divinesolar):

I do understand...

OpenStudy (anonymous):

thank you @DivineSolar

DivineSolar (divinesolar):

Yes sure thing. I hope my worked out version of it explained how to do it :)!